OpenStudy (anonymous):

(3ab)(-5ab) i need to know how to do it

OpenStudy (unklerhaukus):

\[(3ab)(-5ab)=3\times a\times b\times (-5)\times a\times b\]

OpenStudy (unklerhaukus):

\[=-5\times 3\times a\times a \times b \times b\]

OpenStudy (anonymous):

I do not get it

OpenStudy (unklerhaukus):

which bit?

OpenStudy (anonymous):

all of it, i just want to know if the answer is \[-15a^{2}b ^{2}\] ?

OpenStudy (unklerhaukus):

yeah that is right ,

OpenStudy (anonymous):

Yay, thanks

OpenStudy (unklerhaukus):

because \(a\times a=a^2\) , \(b\times b=b^2\)

OpenStudy (anonymous):

Can you help me with this question ... Which of the following is the equation of a line that passes through (-2, -1) and (-4, -3) ? Please

OpenStudy (anonymous):

These are my multiple choice answers

OpenStudy (anonymous):

Nvm . Just can you help me ?